OpenStudy (anonymous):

There are 12 women and 16 men in a room. How many different committees of size 5 can be selected if the committee must have exactly 2 women and 3 men? My answer makes no sense to me. I got 110, 880 but that seems far fetched. I did C(12,2)C(16,3)

OpenStudy (kropot72):

There are 12C2 ways of selecting the two women and 16C3 ways of selecting the two men. Total number of ways of selecting committee is given by 12C2 + 16C3

OpenStudy (anonymous):

So then my answer is actually 560?

OpenStudy (anonymous):

Excuse me- I mean 626? 12!/(12-2)!*2!= 66 16!/(16-3)!*3!=560 66+560= 626?

OpenStudy (kropot72):

Yes, 626 is correct.

OpenStudy (anonymous):

Awesomesauce! Thank you!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (kropot72):

Very sorry, my bad. The total number of ways of forming the committee of 5 is \[12C2\times 16C3 = 66\times 560=36960\]